Use the Principle of Mathematical Induction to prove that (2i+3) = n(n + 4) for all n > 1.
Answer #1

E (21+3) = n(n+4) for all nal (2+3) + (4 +3) +(6+3) + ... + (90+3 ) = n(n+4) 7021 We proof by mathemerical induction Ger nel